# ADPro: a Test-time Adaptive Diffusion Policy for Robot Manipulation via Manifold and Initial Noise Constraints

### v1

@@ 3.3 Adaptive Diffusion Policy with Manifold Constraint
To minimize unnecessary backtracking in the diffusion process (illustrated as Diffuser in Fig. 4), we introduce a manifold constraint into the reverse diffusion policy, ensuring that the denoising updates align more closely with the manifold’s geodesic path. Specifically, we adopt the Gaussian spherical prior (Yang et al. 2024b) to refine the reverse diffusion step size. The Gaussian spherical prior constrains the noise sampling space to a uniform distribution on the hypersphere of radius $\sqrt{d}\sigma$, keeping updated step size within high-confidence intervals. Here, $d$ is the product of the action dimension, the batch size, and the number of actions $n$ to be predicted. As remarked in (Yang et al. 2024b), when $d$ is sufficiently large, a $d$-dimensional isotropy Gaussian distribution $\mathcal{N}(0,\sigma^{2}I)$ is close to a uniform distribution on the hypersphere of radius $\sqrt{d}\sigma$. Then, the refined training-free adaptive diffusion policy is formulated as:
@@
3.4 Initial Noise Constraints by Action Proposal
@@
Given the end-effector $\mathbf{O}_{k,0}$ and a target scene $\mathbf{O}_{k,1}$, the search space for initial actions should be restricted to a specific local region rather than spanning the entire noise space, as is common in existing diffusion policies. Relying on a predefined noise distribution for sampling across the full noise space not only increases the complexity of subsequent denoising but also lowers the manipulation success rate. A more effective way is to propose a reasonable initial action by computing a coarse registration, which simplifies denoising and reduces the number of diffusion steps needed.
Specifically, we employ Fast Global Registration (FGR) algorithm (Zhou, Park, and Koltun 2016) to achieve coarse registration between the observed data $\mathbf{O}_{k,0}$ and $\mathbf{O}_{k,1}$. The resulting output serves as the proposed initial noisy action at step $M$ ($M\leq T$). Compared to the classical Iterative Closest Point (ICP) (Besl and McKay 1992), FGR offers improved efficiency and robustness, particularly when dealing with noisy or large-scale point clouds.
$\mathbf{a}_{k}^{M}=FGR(\mathbf{O}_{k,0},\mathbf{O}_{k,1})$ (10)
Figure 3: Trajectory comparison between the vanilla diffusion policy and our ADPro. We illustrate two distinct tasks: ‘put groceries in cupboard’ and ‘insert peg’. For each task, we visualize the initial state followed by successive observations captured from the wrist-mounted camera after each action. ADPro completes the first task with only three actions, demonstrating both efficiency and accuracy.
3.5 Implementation and Error Analysis of ADP
For the robotic manipulation, ADP can generate control actions and be seamlessly integrated with other DPs in a plug-and-play fashion, without requiring model retraining. Specifically, we first generate noisy initial actions $\mathbf{a}_{k}^{T}$ using Eq. (10), and then iteratively denoise them with Eq. (5), Eq. (7) and Eq. (9) until the final noiseless action $\mathbf{a}_{k}^{0}$ is obtained.
@@ 3.5 Implementation and Error Analysis of ADP
Table 1: The mean success rate on RLBench across three random seeds. ADPro outperforms all prior arts on most tasks by a large margin. Bold indicates the best; underlining shows the suboptimal.
@@
Avg. open slide sweep to put in turn put in place stack sort insert drag place close meat off stack push screw put in Success $\uparrow$ drawer block dustpan safe tap drawer wine cups shape peg stick cups jar grill blocks buttons bulb cupboard C2F-ARM 20.1 20 16 0 12 68 4 8 0 8 4 24 24 0 20 0 72 8 0 PerAct 49.4 88.0 74.0 52.0 86.0 88.0 51.2 44.8 2.4 16.8 5.6 89.6 2.4 55.2 70.4 92.8 0 17.6 28.0 HiveFormer 45 52 64 28 76 80 68 80 0 8 0 76 0 52 100 8 84 8 32 PolarNet 46 84 56 52 84 80 32 40 8 12 5 92 0 36 100 4 96 44 12 RVT 62.9 71.2 81.6 72.0 91.2 93.6 88.0 91.6 26.4 36.0 11.2 99.2 4.0 52.0 88.0 28.8 100 48.0 49.6 Act3D 63.2 78.4 96.0 86.4 95.2 94.4 91.2 59.2 9.6 29.6 24.0 80.8 3.2 96.8 95.2 4.0 93.6 32.8 67.2 Diffuser 81.3 89.6 97.6 84.0 97.6 99.2 95.0 93.6 47.2 44.0 65.6 100.0 24.0 96.0 96.8 66.8 98.4 82.4 85.6 VPDD 56.2 - 70.7 40.0 70.7 88.7 24.0 60.0 56.0 - 73.7 66.7 30.7 37.3 73.3 56.0 58.7 61.3 30.67 Ours 83.9 93.1 100.0 97.0 98.5 99.5 95.2 94.0 56.5 50.0 74.5 100.0 25.4 96.2 95.6 67.3 98.6 83.1 85.9
@@
Figure 4: Comparison on components ($x$, $z$, $r_{1}$, $r_{2}$) of the full action for task ‘sweep to dustpan’. Full action comparisons are available in the supplementary. The horizontal and vertical axes represent action steps and parameter values, respectively. Our ADPro effectively mitigates coordinate and angle backtracking behaviors in the vanilla diffusion policy.
@@
Considering that 3D robot policies have shown superior generalization over 2D policies across varying camera viewpoints and are more effective in handling novel viewpoints during testing (Ke, Gkanatsios, and Fragkiadaki 2024; Ze et al. 2024; Jia et al. 2025), we implement ADP with 3D Diffuser Actor to create ADPro. ADPro’s conditional noise predictor $\varepsilon_{\theta}$ , leverages a pretrained 3D Transformer-based network. To further enhance the model’s generalization across diverse tasks, we incorporate task instructions as textual prompts encoded by the text encoder of the vision-language model CLIP (Radford et al. 2021). The resulting text features are then fused with visual features through a cross-attention mechanism, enabling a more effective integration of semantic and visual information.
Error Analysis.
@@
The robustness of the proposed ADP depends on both the FGR algorithm and the pre-trained diffusion model. When $\mathbf{a}_{k}^{M}$ in Eq. 10 satisfies $\mathbf{a}_{k}^{M}-\mathbf{a}_{k}^{0}||<\delta_{1}$, the noisy action $\hat{\mathbf{a}}_{k}^{M}$ of the pretrained $\varepsilon_{\theta}$ at steps $M$ satisfies $\hat{\mathbf{a}}_{k}^{M}-\mathbf{a}_{k}^{0}||<\delta_{2}$, then, $\mathbf{a}_{k}^{M}-\hat{\mathbf{a}}_{k}^{M}||<\delta_{1}+\delta_{2}$. Furthermore, $\delta_{1}$ is controllable and can be reduced by increasing the number of iterations of the FGR algorithm, which in turn lowers $\delta_{1}$ and reduces the error upper bound. This ensures the robustness of the proposed ADP.
4 Experiments
@@
We designed experiments to answer the following questions: i) Does the proposed diffusion policy improve manipulation performance? ii) How does ADPro perform in terms of generalization and efficiency? iii) Can ADPro be applied to real-world robotic manipulation tasks? We first demonstrate the performance improvements of our proposed ADPro on RLBench (James et al. 2020) and CALVIN (Mees et al. 2022). Then, in Sections 4.3 and 4.4, we validate its enhancements in generalization and efficiency on CALVIN and a real-world dataset. Finally, we conduct ablation studies in Section 4.5 to verify the effectiveness of each core module.
4.1 Baselines
@@
We evaluate the proposed methods with thirteen baselines in four datasets, including C2F-ARM (James et al. 2022), PerAct (Shridhar, Manuelli, and Fox 2023), HiveFormer (Guhur et al. 2023), PolarNet (Chen et al. 2023), RVT (Goyal et al. 2023), Act3D (Gervet et al. 2023), HULC (Mees, Hermann, and Burgard 2022), RoboFlamingo (Li et al. 2024), SuSIE (Black et al. 2024b), GR-1 (Wu et al. 2024), and three DPs: 3D Diffusion Policy (DP3) (Ze et al. 2024), 3D Diffuser Actor (Diffuser) (Ke, Gkanatsios, and Fragkiadaki 2024), VPDD (He et al. 2025). For VPDD, we use the results reported in their official paper. Since no publicly available pretrained model of DP3 for CALVIN exists, we train it on CALVIN to ensure a fair comparison. For the remaining baselines, we adopt the results from 3D Diffuser Actor. Firstly, we leverage benchmarks RLBench and CALVIN to present the effectiveness of the proposed ADP on task success rates and inference speed. Meanwhile, we demonstrate the generalization of our policy in real-world scenes with Franka Panda Robot. We also compare against ablative versions that do not consider manifold or initial noise constraints.
4.2 Performance on RLBench
@@
RLBench is a simulation environment where a Franka Panda Robot is used to manipulate the scene with BiRRT (Kuffner and LaValle 2000) as its motion planner. On RLBench, all methods need to predict the next end-effector keypose. We evaluate ADPro on 18 tasks as  (Ke, Gkanatsios, and Fragkiadaki 2024), each with 2-60 variations. We evaluate policies by task completion success rate, the ratio of execution trajectories that achieve the goal conditions specified in the language instructions (Gervet et al. 2023; Shridhar, Manuelli, and Fox 2023). Tab. 1 presents the evaluation metrics. The baseline performances are reported from Diffuser, and our results are averaged over three random seeds.
@@
As shown in Table 1, our method outperforms the vanilla diffusion policy (Diffuser), with particularly notable improvements on the tasks ‘sweep to dustpan’, ‘sort shape’, and ‘insert peg’. Additionally, we show the robot arm’s Cartesian pose and gripper-gap evolution during a ‘sweep to dustpan’ episode in Fig. 4. From the plots of $r_{2}$ and $r_{3}$, it can be seen that our approach yields more consistent values, and it completes the task using only four actions—one fewer than the five used in the ground-truth demonstration. This reduction in action count stems from our guided diffusion mechanism, which accelerates the grasping process. In Fig. 5, we examine the inverse diffusion process for each action prediction. To enhance readability, we compute the Mean Squared Error (MSE) between the predicted actions and the ground truth, where values closer to zero indicate better predictions. The results demonstrate that incorporating our guidance during generation substantially reduces unnecessary backtracking and improves stability, yielding a much smoother generation trajectory. Notably, for ‘Action 2’ and ‘Action 4’, due to alignment, the initial position of the diffusion process is closer to the ground truth, which facilitates faster convergence toward an accurate prediction and reduces the number of diffusion steps required. In the left panel of Fig. 6, we vary the number of diffusion steps used for generation and evaluate the average success rate on RLBench. We find that our method significantly outperforms the baseline, particularly when using a small number of diffusion steps. Notably, our method achieves a $72\%$ success rate with only 20 steps, whereas the baseline requires at least 50 steps to reach comparable performance. Moreover, our method consistently improves upon the baseline across all diffusion step settings.
Table 2: Zero-shot long-horizon evaluation on CALVIN.
@@
4.3 Generalization on CALVIN
The CALVIN benchmark is built on the PyBullet (Coumans and Bai 2016) simulator and features a Franka Panda robot arm interacting with its scene. It includes 34 tasks across four environments (A, B, C, and D), each equipped with a desk, sliding door, drawer, LED button, lightbulb switch, and three colored blocks. The environments vary in desk texture and object placement. CALVIN provides 24 hours of unstructured play data, with 35% annotated using natural language.
We evaluate models under the zero-shot generalization setup: training on environments A, B, and C, and testing on D. Since CALVIN lacks a motion planner, models must directly predict robot pose trajectories. For a fair comparison, we follow the same evaluation protocol as Diffuser. Each method predicts at most 60 actions per sequence and is evaluated across three random seeds. We report the success rate and the average number of sequential tasks completed. Additionally, we compare DP3 and Diffuser in terms of task completion time (in seconds) and the number of diffusion steps (NDS) required, with results shown in the final column. Other methods are excluded from this comparison, as they do not use diffusion-based policies.
As shown in Tab. 2, our method (last row) consistently outperforms existing approaches on the zero-shot long-horizon prediction task, achieving a roughly 10-point improvement in the success rate of the 5th key pose and producing trajectories with the highest average length. Furthermore, pairwise comparisons between DP3 and DP3+ADP (which integrates our proposed ADP into DP3), and between Diffuser and our method (Diffuser with ADP) highlight the effectiveness of ADP as a plug-and-play module. ADP significantly improves the generalization performance of DPs and boosts generation efficiency, reducing time costs by approximately 25%.

### v2

@@ III-C Manifold-Constrained Denoising
Figure 3: Trajectory comparison between the vanilla diffusion policy and our ADPro. We illustrate two distinct tasks: ‘sort shape’ and ‘insert peg’. For each task, we visualize the initial state followed by successive observations captured from the wrist-mounted camera after each action. ADPro completes the first task with only three actions, demonstrating both efficiency and accuracy.
Spherical manifold
To minimize unnecessary backtracking in the diffusion process (illustrated as Diffuser in Fig. 4), we introduce a manifold constraint into the reverse diffusion policy, ensuring that the denoising updates align more closely with the manifold’s geodesic path. Specifically, we adopt the Gaussian spherical prior [70] to refine the reverse diffusion step size. The Gaussian spherical prior constrains the noise sampling space to a uniform distribution on the hypersphere of radius $\sqrt{d}\sigma$, keeping updated step size within high-confidence intervals. Here, $d$ is the product of the action dimension, the batch size, and the number of actions $n$ to be predicted. As remarked in [70], when $d$ is sufficiently large, a $d$-dimensional isotropic Gaussian distribution $\mathcal{N}(0,\sigma^{2}I)$ is close to a uniform distribution on the hypersphere of radius $\sqrt{d}\sigma$. Then, the refined training-free adaptive diffusion policy is formulated as:
@@
By unifying these constraints under the same principle of manifold-constrained denoising, ADP transforms the reverse diffusion process into a structured trajectory: task manifolds provide directionality, spherical manifolds offer stability, and together they yield more efficient and robust action generation.
III-D Task-Aware Initialization
@@
Given the end-effector $\mathbf{O}^{k,0}$ and a target scene $\mathbf{O}^{k,1}$, the search space for initial actions should be restricted to a specific local region rather than spanning the entire noise space, as is common in existing diffusion policies. Relying on a predefined noise distribution for sampling across the full noise space not only increases the complexity of subsequent denoising but also lowers the manipulation success rate. A more effective way is to propose a reasonable initial action by computing a coarse registration, which simplifies denoising and reduces the number of diffusion steps needed.
To this end, ADP introduces a task-aware initialization module. Rather than sampling from an uninformed prior, we compute a coarse alignment between the gripper and the target scene point clouds. In our implementation, we employ the Fast Global Registration (FGR) algorithm [71] to align the observed data $\mathbf{O}^{k,0}$ with $\mathbf{O}^{k,1}$. FGR outputs a rotation matrix $\mathbf{R}\in\mathrm{SO}(3)$ and a translation vector $\mathbf{v}\in\mathbb{R}^{3}$, which together define the initial noisy action $\mathbf{a}^{k}_{M}$, corresponding to the noisy output at step $M$ ($M\leq T$) of the original diffusion process. Compared with the classical Iterative Closest Point (ICP) algorithm [72], FGR offers greater efficiency and robustness, particularly in handling noisy or large-scale point clouds.
$\mathbf{a}^{k}_{M}=FGR(\mathbf{O}^{k,0},\mathbf{O}^{k,1})$ (10)
In practice, we randomly sample 4096 points from each point cloud and use FGR to align the end-effector point cloud to the scene point cloud. A maximum of 6–10 iterations is sufficient, as only a coarse estimate is required. By initializing the diffusion process in a task-relevant region, we effectively reduce the search space and accelerate convergence. Ablation studies confirm that removing this component significantly degrades both success rate and efficiency.
III-E Unified Adaptive Diffusion Policy and Error Analysis
Bringing these components together, ADP functions as a test-time optimizer for pretrained diffusion policies: (i) initialization places the process near a plausible region, (ii) manifold constraints guide updates along geometry rather than noise, and (iii) the spherical constraint regularizes step magnitudes to prevent instability. In robotic manipulation, ADP generates control actions and integrates seamlessly with existing diffusion policies in a plug-and-play manner, without retraining. Concretely, we first generate noisy initial actions $\mathbf{a}^{k}_{M}$ using Eq. (10), and then iteratively denoise them via Eq. (5), Eq. (7), and Eq. (9) until the final noiseless action $\mathbf{a}^{k}_{0}$ is obtained.
@@ III-E Unified Adaptive Diffusion Policy and Error Analysis
TABLE I: The mean success rate on RLBench across three random seeds. ADPro outperforms all prior arts on most tasks by a large margin. Bold indicates the best; underlining shows the second best.
@@
Avg. open slide sweep to put in turn put in place stack sort insert drag place close meat off stack push screw put in Success $\uparrow$ drawer block dustpan safe tap drawer wine cups shape peg stick cups jar grill blocks buttons bulb cupboard C2F-ARM 20.1 20 16 0 12 68 4 8 0 8 4 24 24 0 20 0 72 8 0 PerAct 49.4 88.0 74.0 52.0 86.0 88.0 51.2 44.8 2.4 16.8 5.6 89.6 2.4 55.2 70.4 26.4 0 17.6 28.0 HiveFormer 45 52 64 28 76 80 68 80 0 8 0 76 0 52 100 8 84 8 32 PolarNet 46 84 56 52 84 80 32 40 8 12 5 92 0 36 100 4 96 44 12 RVT 62.9 71.2 81.6 72.0 91.2 93.6 88.0 91.6 26.4 36.0 11.2 99.2 4.0 52.0 88.0 28.8 100 48.0 49.6 Act3D 63.2 78.4 96.0 86.4 95.2 94.4 91.2 59.2 9.6 29.6 24.0 80.8 3.2 96.8 95.2 4.0 93.6 32.8 67.2 Diffuser 81.3 89.6 97.6 84.0 97.6 99.2 95.0 93.6 47.2 44.0 65.6 100.0 24.0 96.0 96.8 66.8 98.4 82.4 85.6 VPDD 56.2 - 70.7 40.0 70.7 88.7 24.0 60.0 56.0 - 73.7 66.7 30.7 37.3 73.3 56.0 58.7 61.3 30.67 Ours 83.9 93.1 100.0 97.0 98.5 99.5 95.2 94.0 56.5 50.0 74.5 100.0 25.4 96.2 95.6 67.3 98.6 83.1 85.9
(a) Open drawer
(b) Sweep to dustpan
@@
Figure 4: Comparison on all components ($x$, $y$, $z$, $r_{1}$, $r_{2}$, $r_{3}$, $r_{4}$, $w$) of the full action for tasks ‘open drawer’ and ‘sweep to dustpan’. The horizontal and vertical axes represent action steps and parameter values, respectively. Our ADPro effectively mitigates coordinate and angle backtracking behaviors in the vanilla diffusion policy.
@@
Considering that 3D robot policies have shown superior generalization over 2D policies across varying camera viewpoints and are more effective in handling novel viewpoints during testing [29, 26, 32], we implement ADP with 3D Diffuser Actor to create ADPro. ADPro’s conditional noise predictor $\varepsilon_{\theta}$ , leverages a pretrained 3D Transformer-based network. To further enhance the model’s generalization across diverse tasks, we incorporate task instructions as textual prompts encoded by the text encoder of the vision-language model CLIP [73]. The resulting text features are then fused with visual features through a cross-attention mechanism, enabling a more effective integration of semantic and visual information.
@@
Error Analysis. The robustness of the proposed ADP depends on both the FGR algorithm and the pre-trained diffusion model. When $\mathbf{a}^{k}_{M}$ in Eq. 10 satisfies $\mathbf{a}^{k}_{M}-\mathbf{a}^{k}_{0}||<\delta_{1}$, the noisy action $\hat{\mathbf{a}}^{k}_{M}$ of the pretrained $\varepsilon_{\theta}$ at steps $M$ satisfies $\hat{\mathbf{a}}^{k}_{M}-\mathbf{a}_{k}^{0}||<\delta_{2}$, then, $\mathbf{a}^{k}_{M}-\hat{\mathbf{a}}^{k}_{M}||<\delta_{1}+\delta_{2}$. Furthermore, $\delta_{1}$ is controllable and can be reduced by increasing the number of iterations of the FGR algorithm, which in turn lowers $\delta_{1}$ and reduces the error upper bound. This ensures the robustness of the proposed ADP.
IV Experiments
@@
We designed experiments to answer the following questions: i) Does the proposed diffusion policy improve manipulation performance? ii) How does ADPro perform in terms of generalization and efficiency? iii) Can ADPro be applied to real-world robotic manipulation tasks? We first demonstrate the performance improvements of our proposed ADPro on RLBench [74] and CALVIN [75]. Then, in Sections IV-C and IV-D, we validate its enhancements in generalization and efficiency on CALVIN and a real-world dataset. Finally, we conduct ablation studies in Section IV-E to verify the effectiveness of each core module.
IV-A Baselines
@@
We evaluate the proposed methods with thirteen baselines in four datasets, including C2F-ARM [76], PerAct [77], HiveFormer [78], PolarNet [79], RVT [80], Act3D [81], HULC [82], RoboFlamingo [53], SuSIE [83], GR-1 [84], and three DPs: 3D Diffusion Policy (DP3) [26], 3D Diffuser Actor (Diffuser) [29], VPDD [22]. For VPDD, we use the results reported in their official paper. Since no publicly available pretrained model of DP3 for CALVIN exists, we train it on CALVIN to ensure a fair comparison. For the remaining baselines, we adopt the results from 3D Diffuser Actor. Firstly, we leverage benchmarks RLBench and CALVIN to present the effectiveness of the proposed ADP on task success rates and inference speed. Meanwhile, we demonstrate the generalization of our policy in real-world scenes with Franka Panda Robot. We also compare against ablative versions that do not consider manifold or initial noise constraints.
IV-B Performance on RLBench
@@
RLBench is a simulation environment where a Franka Panda Robot is used to manipulate the scene with BiRRT [85] as its motion planner. On RLBench, all methods need to predict the next end-effector keypose. We evaluate ADPro on 18 tasks as  [29], each with 2-60 variations. We evaluate policies by task completion success rate, the ratio of execution trajectories that achieve the goal conditions specified in the language instructions [81, 77]. Tab. I presents the evaluation metrics. The baseline performances are reported from Diffuser, and our results are averaged over three random seeds.
As shown in Table I, our method consistently outperforms the vanilla diffusion policy (Diffuser), with particularly notable gains on the tasks ‘sweep to dustpan’, ‘sort shape’, and ‘insert peg’, achieving an average success rate improvement of 2.6 points. Figure 3 further illustrates these advantages: on ‘sort shape’, ADPro completes the task in three direct actions, whereas Diffuser requires four with detours; on ‘insert peg’, ADPro achieves early alignment and inserts without backtracking, while Diffuser needs an extra alignment action to finish the task. This visual comparison reinforces the efficiency and stability benefits of our approach. Furthermore, Fig. 4 compares all parameters of a 7-DOF robotic arm—including position ($x$, $y$, $z$), rotation ($r_{1}$, $r_{2}$, $r_{3}$, $r_{4}$), and gripper width (ww)—across tasks. Our method converges faster and more accurately to the target grasp configuration, while Diffuser exhibits significant backtracking. These results validate the effectiveness of our guidance mechanism and initialization strategy. Notably, ADPro completes tasks in fewer steps, demonstrating superior efficiency in generating high-quality manipulation trajectories.
@@
In Fig. 5, we examine the inverse diffusion process for each action prediction. To enhance readability, we compute the Mean Squared Error (MSE) between the predicted actions and the ground truth, where values closer to zero indicate better predictions. The results demonstrate that incorporating our guidance during generation substantially reduces unnecessary backtracking and improves stability, yielding a much smoother generation trajectory. Notably, for ‘Action 2’ and ‘Action 4’, due to alignment, the initial position of the diffusion process is closer to the ground truth, which facilitates faster convergence toward an accurate prediction and reduces the number of diffusion steps required. In the left panel of Fig. 6, we vary the number of diffusion steps used for generation and evaluate the average success rate on RLBench. We find that our method significantly outperforms the baseline, particularly when using a small number of diffusion steps. Notably, our method achieves a $72\%$ success rate with only 20 steps, whereas the baseline requires at least 50 steps to reach comparable performance. Moreover, our method consistently improves upon the baseline across all diffusion step settings.
TABLE II: Zero-shot long-horizon evaluation on CALVIN.
@@
IV-C Generalization on CALVIN
The CALVIN benchmark is built on the PyBullet [86] simulator and features a Franka Panda robot arm interacting with its scene. It includes 34 tasks across four environments (A, B, C, and D), each equipped with a desk, sliding door, drawer, LED button, lightbulb switch, and three colored blocks. The environments vary in desk texture and object placement. CALVIN provides 24 hours of unstructured play data, with 35% annotated using natural language.
We evaluate models under the zero-shot generalization setup: training on environments A, B, and C, and testing on D. Since CALVIN lacks a motion planner, models must directly predict robot pose trajectories. For a fair comparison, we follow the same evaluation protocol as Diffuser. Each method predicts at most 60 actions per sequence and is evaluated across three random seeds. We report the success rate and the average number of sequential tasks completed. Additionally, we compare DP3 and Diffuser in task completion time (in seconds) and the number of diffusion steps (NDS) required, with results shown in the final column. Other methods are excluded from this comparison, as they do not use diffusion-based policies.
As shown in Tab. II, our method (last row) consistently outperforms existing approaches on the zero-shot long-horizon prediction task, achieving a roughly 10-point improvement in the success rate of the 5th key pose and producing trajectories with the highest average length. Furthermore, pairwise comparisons between DP3 and DP3+ADP (which integrates our proposed ADP into DP3), and between Diffuser and our method (Diffuser with ADP) highlight the effectiveness of ADP as a plug-and-play module. For DP3+ADP, we only adjust the input format and evaluation protocol to align with the CALVIN benchmark, without altering the core model. Overall, ADP substantially improves the generalization of DPs while also boosting generation efficiency, reducing computational time by approximately 25%.